Hitachi FB-2000A FIB Archive February 3, 2008 – February 7, 2012

Hitachi FB-2000A FIB

Hitachi FB-2000A FIB

February 7, 2012 I have an experiment running in the FIB and I can’t get back to it until Friday. Please let me know if you need the FIB tomorrow and I’ll remove the sample myself.

Thanks, Owen

September 29, 2011 I will be taking the FIB off line today until next week. I’ll let you know when it can be used again

July 11, 2011 We are transitioning over to a server-based logbook system. The Apple computers are being replaced with a very simple client box, however the USER INTERFACE IS UNCHANGED. The FIB system is first to be replaced.

Let me know right away if you observe an error of some sort or become confused.

June 23, 2011 The FIB is back online and available for your use.

June 21, 2011 FIB Down: The FIB has lens stability problems. I’m in touch with the Hitachi engineer. Look for update messages as we chase down the problem.

September 25, 2010 The new source we installed in early August is still stabilizing. You may notice a few different examples of this. First, the Vext and Iext may be out of wack. You should see a Iext of ~3.2uA and a corresponding Vext of 8.5kV. If not, make a adjustment to the Iext slider bar in the LMIS menu. The second thing that you may notice is instability over an hour or more of use. As soon as you see instability, reform the beam in the LMIS menu.

Let me know if you experience these problems.

September 7, 2010 We will likely be taking on a special FIB/TEM project late this month or early October. I’ll need 3 continuous days of FIB and 2 days of TEM.

Please let me know ASAP if this will cause you problems.

July 30, 2010 The new tip is working well but maybe not completely stable yet. You may have to flash the tip if you see a current drop. Probably not necessary more than once every 3-5 hours.

July 23, 2010 I will install the SEM stage in the FIB this afternoon for a Monday project. I’ll swap the TEM stage back in after we are done and the scope will be ready to use Tuesday AM.

July 6, 2010 I have ordered the repair parts and Hitachi will begin installation next week. The FIB should be back online ~ July 15 or 16.

July 2, 2010 The FIB is down with an expired emitter. I will let you know the status of the repair but I cannot imagine that it will be back online before mid to late July.

June 1, 2010 The SEM tage is installed on the FIB now. I’ll need 1 day notice if anyone needs the TEM stage back in this week. We have external clients coming in next week to use the FIB Mon – Wed, so it will be unavailable till late next week.

Plan your use accordinly.

May 12, 2010 The FIB will be down for annual maintenance next week, Monday through Wed, maybe into Thursday.

March 29, 2010 The FIB trackball is working intermittently. I’m ordering a new one but it may not be here until next week.

February 22, 2010 I still have the SEM stage installed on the FIB. Let me know if you need the other stage this week.

February 16, 2010 I have a TEM sample prep project to do Thursday. Wed evening, I will install the SEM stage and let it pump overnight. I’ll pull the SEM stage when we are done Thursday afternoon. The FIB, with TEM stage, will be ready for use Friday morning.

December 21, 2009 I have installed a software operated NPGS enable switch to replace the blue manual switch we use now. The new switch is controlled by 2 radio buttons in the LH NPGS control panel, labelled “FIB” and “NPGS”. “FIB” corresponds to the “SEM” switch setting we use now.

June 18, 2009 We found a small hole in a part that carries cooling water. It will have to be replaced. At this time, we expect the Hitachi engineer to return next week or the week after to make the repair. The repair might take 2 days. The system is running and can be used. PLan accordingly.

April 28, 2009 The FIB is back online. The problem was a bad mouse.

April 27, 2009 The software will not load after booting the computer. Waiting for Hitachi to respond.

February 6, 2009 The NPGS computer is back and working fine. Ready to use.

February 4, 2009 The NPGS computer is not working. I’ll bring it in for repair today.

February 2, 2009 Be aware that I’ve changed the deposition gun position slightly.

January 28, 2009 The new NPGS system is installed. I have adjusted the calibration and I think it is very close. Let me know if this is not true. The software is a bit different but close enough to be familiar.

January 23, 2009 Check out this website.

This is a great website for FIB related info with sections on: European Focused Ion Beam Users Group

EFUG presentations sorted by topic (since 2000)

December 5, 2008 Water filter dirty, filter canister begins leaking. Replaced filter and canister, eliminated the in-line pressure regulator. Water pressure surging, tank level low. FIB is back on-line

November 4, 2008 There was a report of short-term beam current fluctuations and out-of-range beam currents. I checked it out today and here’s the results.

Short term beam current changes. FIB beam stability is specified at 5% per hour. During my tests today I did not observe changes outside that range.

Out of range beam currents. I aligned the FIB column today and all beam currents are within specs. Note that the beam currents will change over a longer period (months). If your work is very sensitive to longer term beam current changes you need to keep a log so we can decide when to align the column. It takes a few hours to align the column and I don’t mind doing it as often as needed.

A final note. You have got to let me know if you are having trouble achieving an experimental result that may be related to the instrument.

November 3, 2008 Reports of fluctuating beam current. Extraction current is stable. I’ll realign the scope tomorrow.

October 31, 2008 Re-plumbed the vacuum lines to steel on the service alley side of wall. All seems fine.

Oct 24, 2008 The recirculator is not working. I replaced the pump, cleaned all the filters, changed the water. All is well now. I did not test the beam but I don’t expect problems since the ion pumps were on during the repair

October 10. 2008 The FIB depo gun is working properly and is height adjusted. Do be sure to check the height adjustment anyway. Also, Please use M0-100 for depo work, NOT M0-50.

October 8, 2008 I have replenished the depo gun W supply and aligned the nozzle to be very close to the specimen surface.

August 22, 2008 The W supply in the depo gun may be depleted. I’ll check it today.

August 8, 2008 If you want to use the depo gun, please mount your sample in the holder with back handle. I have set the depo gun height for that holder. Read on to learn more. Once set, the depo gun position will work with a one particular specimen holder. If the other holder is used for depo, it is likely the gun will hit the specimen.

July 31, 2008 I am planning to swap the SEM stage in for one day tomorrow.

July 18, 2008 The depo gun is pumped out and the beam is stable. FIB is up and running and ready to use.

July 17, 2008 We have completed the swap back to the TEM stage. The WCO compound in the depo gun adsorbs moisture when exposed to atmosphere (during chamber vents to swap stages) and has to be pumped out. We are continuing to pump moisture out of depo gun. That may take remainder of today and even into tonight.

July 8, 2008 The ion beam has been stable for 36 hours. Please begin using the scope. Let me know if you observe any problems.

June 30, 2008 Sorry but the beam stability was short lived. Don’t use the FIB. I’ll keep you updates.

June 30, 2008 The ion beam is stable this afternoon. You may begin using the FIB. Please let me know if anything changes.

June 16, 2008 The ion beam is unstable. I am working with Hitachi to resolve this problem. In the mean time please do not use the FIB.

June 17, 2008 I’m having trouble with the new NPGS install so I will temporarily reinstall the old system. You may begin FIB use again tomorrow morning.

June 6, 2008 I am completing installation of the NPGS system. This process will involve calibrating the scale of milled dimensions to actual dimensions. This process requires using the FESEM to verify actual dimensions. After the installation is complete you all will have to become familiar with the new user interface and CAD program. Expect to spend more time for your upcoming work.

June 3, 2008 I am installing the new NPGS system on the FIB. I started this afternoon and should be done by weeks end, if not sooner. I have placed the “Do Not Use…” sign on the door, but it is okay to use it if you will be using the on-board software.

March 28, 2008 The Hitachi engineers will perform routine maintenance on both the FESEM and FIB during the week of April 14th. The scopes will be unavailable after noon on the 14th and will be back on line by noon April 18th. Plan your work accordingly.

February 3, 2008 FIB, TEM stage is reinstalled. scope is working fine, no reports of beam instability.

NPGS. we have 2 new, latest version NPGS system. One will be installed on the 6400 for e-beam litho. The other will replace the one on the FIB. This installation will result in some downtime for NPGS users. I HIGHLY recommend that you back up your patterns to a disk before this switch. I will be scheduling this install with FIB users.

Routine maintenance – The FIB, FESEM and 6400 are all due for factory maintenance. In each case the scope will be down most of a week. I’ll alert you to this schedule.